### v2

# Large Language Models as Span Annotators

Abstract
Span annotation is the task of localizing and classifying text spans according to custom guidelines. Annotated spans can be used to analyze and evaluate high-quality texts for which single-score metrics fail to provide actionable feedback. Until recently, span annotation was limited to human annotators or fine-tuned models. In this study, we show that large language models (LLMs) can serve as flexible and cost-effective span annotation backbones. To demonstrate their utility, we compare LLMs to skilled human annotators on three diverse span annotation tasks: evaluating data-to-text generation, identifying translation errors, and detecting propaganda techniques. We demonstrate that LLMs achieve inter-annotator agreement (IAA) comparable to human annotators at a fraction of a cost per output annotation. We also manually analyze model outputs, finding that LLMs make errors at a similar rate to human annotators. We release the dataset of more than 40k model and human annotations for further research. [^1] Project website: https://llm-span-annotators.github.io
1 Introduction
Fine-grained aspects of texts, such as faithfulness or coherence, depend on local lexical choices. In order to reflect these aspects in quality judgments of texts, we need analysis techniques that provide the appropriate amount of detail. However, most automatic evaluation metrics for natural language generation (NLG) are based on direct assessment, producing holistic scores for each evaluated aspect Gkatzia & Mahamood (2015); Sai et al. (2023); Schmidtová et al. (2024). Although numerical values make it easy to rank systems, these metrics are too simplistic and susceptible to biases or miscalibration of the underlying models Gehrmann et al. (2023); Gao et al. (2024); Liu et al. (2024); Wang et al. (2024).
The subject of our study, span annotation (Figure 1), offers an alternative approach. Instead of assigning a single score for each evaluated aspect, the goal of span annotation is to localize text spans and classify them according to task-specific guidelines. In contrast to numerical rating, annotations are aligned to the evaluated text, which makes them more explainable and actionable. The annotations can also be examined post hoc, which brings more accountability into the evaluation process.
Despite its advantages, span annotation has not yet been widely applied in automatic NLG evaluation. The method traditionally required human annotators, making it costly and difficult to scale Da San Martino et al. (2019); Thomson & Reiter (2020); Popovic (2020); Kocmi et al. (2024c). The situation may change with the upcoming LLM-as-a-judge paradigm, in which LLMs follow instructions for the evaluation process similarly to human annotators. However, current LLM-as-a-judge applications still focus mainly on document-level classification or numerical ratings on the Likert scale Zheng et al. (2023); Gu et al. (2024).
In this work, we investigate whether LLMs can be applied as a substitute for human annotators on span annotation tasks. First, we focus on how to produce span annotations with LLMs. We discuss input tagging methods and experiment with prompting techniques, looking for a setup that is robust and widely applicable. Next, we compare LLMs with skilled human annotators in terms of IAA. Where possible, we contrast the IAA between LLMs and human annotators with the upper bound of the IAA between the annotators themselves. Finally, we manually analyze a subset of LLM and human annotations to assess the validity of the annotations and their respective explanations.
For our experiments, we select three diverse span annotation tasks (cf. Section 3.1): evaluating data-to-text generation Thomson & Reiter (2020), identifying errors in machine translation Kocmi et al. (2024a), and detecting propaganda techniques in human-written texts Da San Martino et al. (2019). Note that unlike established span annotation tasks such as named entity recognition Tjong Kim Sang & De Meulder (2003); Wang et al. (2025a), there are no general purpose fine-tuned models for these tasks since annotation guidelines are specified ad hoc, making their automation difficult.
We base our span annotators on open and proprietary state-of-the-art LLMs, including recent reasoning LLMs that leverage test-time compute for better performance on challenging tasks Muennighoff et al. (2025); Marjanović et al. (2025).
Our contributions are as follows:
  1. 1.
    We demonstrate that zero-shot prompting in combination with structured outputs can be a robust way to set up LLMs as span annotators, given that the model is provided detailed annotation guidelines (Sections 3.2 and 5.1)
  2. 2.
    Using a combination of specialized IAA metrics and manual analysis, we show that performance of LLMs is close to skilled human annotators, in some scenarios leveling or surpassing the performance of qualified crowdworkers (Sections 5.2 and 5.3).
  3. 3.
    We release a dataset of more than 40k human and model annotations for all our tasks and setups, including reasoning traces from reasoning LLMs and annotations collected from skilled crowdworkers.
Figure 1: Examples of span annotation tasks that we automate with LLMs. We unify the setup for evaluation tasks (D2T-Eval, MT-Eval) and text analysis tasks (Propaganda).
2 Related Work
LLMs for NLG Evaluation.
Automatic NLG metrics traditionally assess text quality by measuring similarity to human-written reference texts Sai et al. (2023); Schmidtová et al. (2024). As such, they are unable to quantify more fine-grained aspects (Gehrmann et al., 2023; Freitag et al., 2021) and do not correlate well with human judgments Novikova et al. (2017); Reiter (2018). With the emerging LLM-as-a-judge paradigm (Gu et al., 2024), LLMs have been applied as evaluators across diverse tasks, either with simple numeric scoring (Bavaresco et al., 2024; Liu et al., 2023; Sottana et al., 2023; Leiter et al., 2023; Chiang & Lee, 2023) or with detailed feedback (Li et al., 2024; Kim et al., 2024a; b; Xu et al., 2023). However, the scores produced by LLMs still miss fine-grained text aspects and are influenced by LLM biases Stureborg et al. (2024); Koo et al. (2024); Wang et al. (2024).
Span Annotation Protocol.
In machine translation (MT), span annotation is a long-standing component of protocols such as MQM or ESA (Lommel et al., 2014; Mariana, 2014; Popovic, 2020; Kocmi et al., 2024c), where annotators mark erroneous spans in translations. In data-to-text (D2T) generation, span annotation is applied by Thomson & Reiter (2020), who introduce a span-based evaluation protocol for annotation of generated basketball reports. Span annotation is also used to judge intrinsic text qualities, such as coherence or use of rhetorical devices, in tasks such as propaganda detection Da San Martino et al. (2019) and text summarization Subbiah et al. (2024). Note that unlike our work, these works perform span annotation only using human annotators.
Automating Span Annotation.
For all the aforementioned tasks, first attempts at automating span annotation were based on fine-tuned pre-trained encoder models. That includes MT (Guerreiro et al., 2024), D2T generation Kasner et al. (2021), text summarization Goyal et al. (2022), and propaganda detection Martino et al. (2020); Goffredo et al. (2023); Piskorski et al. (2023). Automating span annotation with LLMs is more flexible and benefits from increasing LLM capabilities. We build on recent work that applies LLMs as a task-specific evaluation tool (Kocmi & Federmann, 2023; Fernandes et al., 2023; Hasanain et al., 2024; Kasner & Dušek, 2024; Chang et al., 2024; Zouhar et al., 2025; Kartác et al., 2025; Ramponi et al., 2025). Our work is the first that unifies diverse span annotation tasks within a single framework and systematically compares the performance of LLMs to human annotators.
See more discussion on related work in Appendix F.
3 Automating Span Annotation with LLMs
In this section, we first formally introduce the span annotation process in Section 3.1. Next, we discuss how to automate the process with LLMs in Section 3.2 and how to evaluate the quality of span annotations in Section 3.3.
3.1 Span Annotation: Task Definition
The goal of span annotation is to annotate a text sequence $Y=\langle y_{1},\ldots,y_{n}\rangle$ given:
  • the set of categories $C=\{c_{1},\ldots,c_{k}\}$,
  • the annotation guidelines $\mathcal{G}$ (including category definitions and other instructions),
  • the source $X$ (such as the translation source; empty if we are annotating only intrinsic text aspects).
The goal is to produce a set of annotations $A=\{a_{1},\ldots,a_{m}\}$, where each annotation $a_{i}$ is a tuple $\langle s_{i},e_{i},c_{i},r_{i}\rangle$:
  • $s_{i},e_{i}\in\{1,\ldots,n\},s_{i}<e_{i}$ are the start and end indices of the annotated span,
  • $c_{i}\in C$ is the assigned annotation category,
  • $r_{i}$ is an explanation for the annotation (optional). [^2] We collect the explanations from LLMs, but not from human annotators due to time considerations.
3.2 Automating Span Annotation with LLMs
To automate span annotation, we collect the annotations $A$ for the given input $\langle Y,C,\mathcal{G},X\rangle$ from an LLM:
$A=\textsc{LLM}(\text{prompt}(Y,C,\mathcal{G},X)).$ (1)
How to Identify the Spans?
An annotation needs to contain the start and end indices of the span. For encoder models, the task would be trivially solved by classifying tokens on top of the encoder’s hidden states. However, to the best of our knowledge, there is no established way of tagging the input text with decoder-only models. Of the methods from previous work, we discarded surrounding spans with special XML-like tokens Wang et al. (2025b); Treviso et al. (2024). This method requires the LLM to generate full copy of the input text and does not offer a fallback if the model fails to do so. We also rule out listing start and end indices of spans Hasanain et al. (2024); Ramponi et al. (2025), as Hasanain et al. note that the model frequently generates indices that do not match the corresponding spans in a paragraph. Instead, we adopted the method of listing textual content of the spans Kocmi & Federmann (2023); Zouhar et al. (2025); Kasner & Dušek (2024). Here, the start and end indices of the span are found heuristically by string matching of the span’s textual content. With this approach, we do not need to modify the input and failing to match a single annotation does not influence other matches. [^3] We do note that this approach can misalign the model output by assigning the annotation to the first match, but our preliminary experiments showed that the collisions are rare.
Parsing Model Output
To obtain the annotations, we request the list of annotations in JSON format, following the setup of Kasner & Dušek (2024). This setup allows us to use constrained decoding with a defined JSON scheme, ensuring that the output is syntactically valid. We require each annotation to contain the fields reason (the explanation $r_{i}$), text (the textual content of the span), and type (the integer index of the error category $c_{i}$). [^4] Given the considerations of Castillo (2024), we ensure that the reason field has precedence.
3.3 Evaluating Span Annotations
To compare annotations automatically, we need a notion of similarity between two sets of annotations $\mathcal{A}=\{A_{1},A_{2},\ldots,A_{|Y|}\}$ and $\hat{\mathcal{A}}=\{\hat{A}_{1},\hat{A}_{2},\ldots,\hat{A}_{|Y|}\}$ over a set of texts $\mathcal{Y}=\{Y_{1},Y_{2},\ldots,Y_{|Y|}\}$. We consider three similarity metrics:
Pearson correlation $\rho$ over counts.
This metric simply compares how many spans were annotated for each example:
$\mathrm{CountCorrelation}(\mathcal{A},\hat{\mathcal{A}})=\mathrm{Pearson}(\langle|A_{Y}|,|\hat{A}_{Y}|\rangle_{Y\in\mathcal{Y}})$ (2)
This metric serves as a sanity check: a low value would suggest that an annotator either skips examples or over-annotates, indicating unclear annotation guidelines.
Precision, Recall, and $\mathbf{F_{1}}$.
To quantify the degree of alignment between individual annotations, we compute precision, recall, and $F_{1}$ as defined in Da San Martino et al. (2019). These measures are on matching annotations, adjusted to give partial credit to imperfect matches: $\text{Precision}(A_{Y},\hat{A}_{Y})=\frac{1}{|A_{Y}|}\sum_{a\in A_{Y}}\frac{|a\cap\hat{a}|}{|a|},$ (3) $\text{Recall}(A_{Y},\hat{A}_{Y})=\frac{1}{|\hat{A}_{Y}|}\sum_{\hat{a}\in\hat{A}_{Y}}\frac{|a\cap\hat{a}|}{|\hat{a}|},$ (4) where $a\cap\hat{a}$ is the character overlap between two annotation spans and $|a|=e-s+1$ is the length of the annotation span in characters (see Section 3.1). Subsequently, we compute the $F_{1}$-score as the harmonic mean of precision and recall.
For each of the metrics, we consider soft and hard variants. The hard variant only considers overlaps where the span category matches, while the soft variant disregards the categories. We consider the hard variant to be the default. In addition, we report the difference $F_{1}\Delta=F_{1}(\text{soft})-F_{1}(\text{hard})$.
Gamma $\bm{\gamma}$.
The $F_{1}$ score is sensitive to varying span granularities and does not consider near-matches with no overlap or agreement by chance. To this end, we follow Da San Martino et al. (2019) and Hasanain et al. (2024) in using the $\gamma$ score (Mathet et al., 2015) as a complementary metric. In short, the metric builds the best possible alignment between the sets of annotations $A_{Y}$ and $\hat{A}_{Y}$ and computes the “disorder” of this alignment based on the positional and categorical dissimilarities of aligned annotations. The score ranges from negative infinity to 1, where 1 is achieved when the annotations are perfectly aligned. The $\gamma$ score extends Krippendorff’s $\alpha$ (Krippendorff, 1980), another popular metric, by computing the category-aware span alignments. We use the implementation of Titeux & Riad (2021).
$S_{\emptyset}$ score.
For an output $y$, one or both annotation sets $A_{Y},\hat{A}_{Y}$ can be empty. Having these cases is desirable: e.g. if the goal is to annotate errors in an output that is entirely correct. However, neither of the introduced scores accounts for these cases. [^5] The $\gamma$ score is undefined for less than two non-empty annotation sets: these examples thus need to be skipped during the $\gamma$ computation. The F1 score does not account for examples with no annotations. To compensate for that, we introduce a score $S_{\emptyset}$ that is computed for examples with less than two nonempty sets of annotations as $S_{\emptyset}=1/(1+|A|)$, where:
$|A|=\begin{cases}|A_{Y}|&\text{if }|\hat{A}_{Y}|=0,\\ |\hat{A}_{Y}|&\text{otherwise}.\end{cases}$ (5)
The score is equal to 1 for the cases where no annotator produced any annotation (i.e., a perfect match) and decreases proportionally to the number of annotations from the annotator that provided a non-zero number of annotations.
4 Experiments
Task # Cat. # Texts Avg. Len Novel Data
D2T-Eval 6 1,296 118/715
MT-Eval 2 2,854 26/185
Propaganda 18 100 914/4,659
Table 1: Overview of span annotation tasks used in our experiments. # Cat. denotes the number of categories used in the task (see Appendix C for their listings), # Texts the number of texts annotated, and Avg. Len the average number of words/characters in the output.
We cover three span annotation tasks of different quantitative and qualitative aspects: evaluating data-to-text generation (D2T-Eval; Section 4.1), identifying errors in machine translation (MT-Eval; Section 4.2), and detecting propaganda techniques (Propaganda; Section 4.3). In Table 1, we provide a quantitative overview of our tasks. [^6] See also Figure 1 and Appendix G for the visualization of examples of our tasks.
4.1 Task: Evaluation of Data-to-text Generation
In D2T-Eval, we focus on evaluating the semantic accuracy and stylistic aspects of texts generated from structured data Celikyilmaz et al. (2020); Sharma et al. (2022). The input is the text $Y$ generated from the structured data $X$.
Subtasks and input data.
We select three D2T generation subtasks: generating five-day weather forecasts (openweather), generating product descriptions (gsmarena), and generating reports of soccer games (football). We use the approach of Kasner & Dušek (2024) to download novel structured data from publicly available APIs to make sure that they are unseen for all LLMs (cf. Section 4.4 on data contamination). We provide more details on the data collection process in Section B.1.
Collecting model outputs.
We generate outputs for the structured inputs using LLMs. Specifically, we selected two state-of-the-art models, Llama 3.3 70B Grattafiori et al. (2024) and GPT-4o Hurst et al. (2024), and two smaller models, Gemma 2 2B Team et al. (2024) and Phi 3.5 3.8B Abdin et al. (2024). We prompt the models in a zero-shot setting, asking them to generate a summary of the given data using approximately five sentences. [^7] See Sections A.1 and A.2 for details on the models and Appendix D for the prompts.
Data splits.
For further experiments, we divide the data into multiple subsets:
  • $\mathcal{D}_{\texttt{test}}$ (1200 outputs) – the subset for LLM evaluation, annotated with crowdworkers,
  • $\mathcal{D}_{\texttt{dev}}$ (84 outputs) – the subset for the study of prompt variants, annotated internally,
  • $\mathcal{D}_{\texttt{iaa}}$ (12 outputs) – the control subset for human crowdworkers, annotated internally.
Crowdsourcing annotations
We gather span annotations for $\mathcal{D}_{\texttt{test}}$ with crowdworkers from Prolific.com. We apply best practices for gathering human annotations, including an iterative process of refining annotation guidelines and preselecting best-performing annotators using a qualification task Tseng et al. (2020); Iskender et al. (2020); Huang et al. (2023); Zhang et al. (2023). Of the 230 annotators that participated in the qualification task, we selected 50 best annotators to annotate our data. We provide more details on the process of collecting human annotations in Section B.2.
4.2 Task: Identifying Errors in Machine Translation
For MT-Eval, we use the dataset of system outputs from the WMT 2024 general shared task (Kocmi et al., 2024b). The system outputs were annotated with the Error Span Annotation (ESA) protocol (Kocmi et al., 2024c) by professional translators. We focus on character-level span annotations of major and minor translation errors provided in the dataset. The inputs $X$ in MT-Eval are the source texts used to generate the translation $Y$ [^8] Note unlike for the other tasks, the annotations in MT-Eval cannot overlap and need not to be aligned with word boundaries.
We use data for the three textual domains present in the WMT 2024 shared task: news, literary, and social. We focus on data translated from English into other languages, including Chinese, Czech, German, Hindi, Icelandic, Japanese, Russian, Spanish, Ukrainian.
The original dataset has nearly 50k model outputs, making it too costly to run the full evaluation campaign. Therefore, we used all available system outputs for 10 sampled inputs for each language, giving us 2,854 outputs that we use for the experiments.
4.3 Task: Propaganda Technique Detection
For the Propaganda task, we use the dataset of Da San Martino et al. (2019). The dataset contains news collected from mostly propagandistic on-line sources. The token-level annotations in the dataset created by expert annotators cover 18 categories of logical fallacies and persuasion techniques. We use the test split for our experiments. Note that there are no inputs $X$ for the task, as all the annotated categories are intrinsic to the evaluated text $Y$.
4.4 Mitigating Data Contamination
We took preventive measures to ensure that our results are not affected by data contamination Balloccu et al. (2024); Dong et al. (2024); Jiang et al. (2024b):
  • For D2T-Eval, we follow the approach of Kasner & Dušek (2024) to download a new set of structured data from publicly available sources and generate new outputs.
  • For MT-Eval, we use the WMT 2024 dataset that was released at the end of 2024, which is after the knowledge cut-off date of the LLMs we are using.
  • For Propaganda, the dataset is publicly available since 2019, but due to its format cannot be easily memorized. [^9] The model would need to separately match the annotations for numerical token ids with the text.
4.5 Collecting LLM annotations
Models
For our span annotator LLMs, we use a mix of open and proprietary state-of-the-art models: (1) instruction-tuned models Llama 3.3 70B Grattafiori et al. (2024), GPT-4o Hurst et al. (2024), Claude 3.7 Sonnet Anthropic (2025), and (2) reasoning models: DeepSeek-R1 70B DeepSeek-AI et al. (2025), o3-mini OpenAI (2025) and Gemini 2.0 Flash Thinking Deepmind (2025). We run the open models (Llama 3.3, DeepSeek-R1) locally in 4-bit quantization using ollama.
Prompts
We define several prompt variants for our experiments. $\mathcal{P}_{\texttt{base}}$ is the base prompt that includes the guidelines $\mathcal{G}$ as given to human annotators. By extending $\mathcal{P}_{\texttt{base}}$, we implement few-shot ($\mathcal{P}_{\texttt{5shot}}$) and chain-of-thought ($\mathcal{P}_{\texttt{cot}}$) prompts. We also ablate $\mathcal{P}_{\texttt{base}}$ by removing extended guidelines ($\mathcal{P}_{\texttt{noguide}}$) and not asking for explanations ($\mathcal{P}_{\texttt{noreason}}$). The full prompts can be found in Appendix D.
Annotation Tool
We use factgenie Kasner et al. (2024), a tool that supports collecting span annotations from humans via a web interface and from LLMs via API calls.
5 Results
In this section, we describe the results of our experiments on span annotation with LLMs. We first investigate the effect of prompting techniques in Section 5.1. Next, we evaluate the LLM annotations using automatic metrics (Section 5.2) and manually analyze the errors in the model outputs (Section 5.3).
5.1 Prompting Techniques
We perform a small-scale study on $\mathcal{D}_{\texttt{dev}}$ using open models (Llama 3.3 and DeepSeek-R1) to study the differences between prompting techniques. The results are shown in Table 2.
Llama 3.3 DeepSeek-R1
Prompt $F_{1}$ $\gamma$a/o $F_{1}$ $\gamma$a/o
$\mathcal{P}_{\texttt{base}}$ 0.20 0.13 2.4 0.25 0.20 1.0
$\mathcal{P}_{\texttt{noguide}}$ 0.11 0.08 3.4 0.20 0.16 1.6
$\mathcal{P}_{\texttt{noreason}}$ 0.22 0.13 2.2 0.24 0.18 1.1
$\mathcal{P}_{\texttt{cot}}$ 0.09 0.10 0.8 0.24 0.19 1.1
$\mathcal{P}_{\texttt{5shot}}$ 0.25 0.18 2.5 0.21 0.16 1.4
Table 2: Comparison of prompting techniques on the $\mathcal{D}_{\texttt{dev}}$ (#a/o is the average number of annotations per output).
Including detailed guidelines seems beneficial: omitting the guidelines ($\mathcal{P}_{\texttt{noguide}}$) lowers the performance of both models. In contrast, not letting the model explain the annotation ($\mathcal{P}_{\texttt{noreason}}$) does not have a substantial effect. For Llama 3.3, the chain-of-thought (CoT) prompting ($\mathcal{P}_{\texttt{cot}}$) makes it produce less annotations per example than the base variant (0.8 vs. 2.4), leading to lower F1 and $\gamma$ scores. [^10] The model with $\mathcal{P}_{\texttt{cot}}$ tends to “overthink” the annotations, deciding not to annotate cases of errors against which it can find some arguments. We provide an example of this behavior in Table 27. Few-shot prompting ($\mathcal{P}_{\texttt{5shot}}$) brings ambivalent results, increasing Llama 3.3 scores but doing the opposite for DeepSeek-R1.111111This observation is aligned with DeepSeek-AI et al. (2025), who note that few-shot prompting consistently degrades the performance of DeepSeek-R1. Given these considerations, we decided to use $\mathcal{P}_{\texttt{base}}$ for our main experiments.
5.2 LLM vs. Human Annotations
Figure 2: Comparison of LLMs using $\mathcal{P}_{\texttt{base}}$ and human annotators. The stripe pattern denotes proprietary models, darker shades denote reasoning models. For $F_{1}$ score, the grey bar denotes its soft variant. The dashed horizontal lines for D2T-Eval denote agreement between our human annotators. Detailed results are included in Tables 8, 16 and 20.
Next, we compute IAA between LLM and human annotations using the metrics described in Section 3.3. The overall results for all tasks are given in Figure 2. We also provide detailed results for individual tasks in Appendix E.
Figure 3: Confusion matrix for D2T-Eval (Contradictory, Not checkable, Misleading, Incoherent, Repetitive, Other, cf. Table 5).
Model football gsmarena openweath.
$F_{1}$ $\gamma$ $F_{1}$ $\gamma$ $F_{1}$ $\gamma$
Llama 3.3 0.29 0.20 0.11 0.08 0.12 0.05
GPT-4o 0.30 0.19 0.09 0.06 0.18 0.12
Claude 3.7 0.38 0.25 0.16 0.14 0.27 0.28
DeepS. R1 0.34 0.28 0.12 0.16 0.18 0.12
o3-mini 0.42 0.33 0.15 0.19 0.29 0.31
Gem. 2-FT 0.38 0.31 0.16 0.16 0.22 0.24
Table 3: Performance of models by domain (football, gsmarena, openweather) in D2T-Eval: $F_{1}$ score (hard), $\gamma$ score. The best results are bolded.
Reasoning models outperform instruction-tuned models.
DeepSeek-R1 generally outperforms the same-sized Llama 3.3. The superiority of DeepSeek-R1 is most pronounced on D2T-Eval ($F_{1}$-score of 0.23 vs. 0.18, $\gamma$ score of 0.19 vs. 0.11). The same observation applies to OpenAI models, where o3-mini outperforms GPT-4o despite o3-mini’s price per token being approximately 2x lower. A notable exception to this trend is the Claude 3.7 Sonnet, which is mostly on par with o3-mini and excels on MT-Eval. However, its price per token is approximately 3x higher (cf. Table 4 for model costs).
LLMs reach human IAA on D2T-Eval, Propaganda is harder.
For D2T-Eval, we can compare model results with an average IAA on examples annotated by two human annotators (cf. Section B.2). Here, o3-mini, Claude 3.7 and Gemini 2.0 mostly reach or surpass human agreement. For Propaganda, the upper bound of IAA is the result of Da San Martino et al. (2019), who report $\gamma=0.31$ for annotators before consolidation. This score is substantially higher (the best LLM score being $\gamma=0.16$ for the Claude 3.7 Sonnet). However, this task used expert annotators and has the largest number of categories. The latter manifests itself in the large difference between the soft and hard $F_{1}$ scores. [^12] The WMT24 dataset for MT-Eval contains examples annotated with a pair of annotators, but these examples form only a small fraction and exhibit high variance between language pairs. Therefore, we decided to omit the comparison with human IAA in MT-Eval.
Models confuse related categories.
Confusion matrices (cf. Figure 3 and Section E.8) suggest that the models tend mainly to confuse related categories such as Contradictory, Not checkable and Misleading for D2T-Eval, which may be related to ambiguity or subjective understanding of category definitions. The models also use a less diverse distribution of categories, e.g., resorting to the Contradictory category more often than human annotators.
LLMs are more efficient than human annotators.
For D2T-Eval, crowdsourced annotation for 1k outputs costs approximately \$500, while annotating the same amount of outputs with the high-performance model o3-mini LLM costs \$3.60 (cf. Table 4). The crowdworkers take 129.1 seconds per output on average, which is better than DeepSeek-R1 on our local infrastructure, but order of magnitude slower than API-based models. [^13] Also note that unlike LLMs, the crowdworkers are not required to produce explanations for the annotations. This suggests that LLMs can be an effective alternative to human annotators both in terms of costs and time.
Metric crowdw. Llama 3.3 DeepSeek-R1 Claude 3.7 Sonnet o3-mini
Cost per 1k outputs (\$) 500 - - 10.5 3.6
Time per output (s) 129.1 21.6 227.5 9.0 21.8
Table 4: Estimate of costs and time requirements on D2T-Eval: crowdworkers on Prolific, open models (Llama 3.3, DeepSeek-R1), and proprietary models (Claude 3.7 Sonnet, o3-mini). Note that we omit the infrastructure and energy costs for the models deployed locally.
5.3 Manual Analysis of LLM Annotations
We manually analyzed the quality of LLM annotations on 216 samples from D2T-Eval and Propaganda. For each model, we sampled three annotations per category in D2T-Eval and one annotation per category in Propaganda. We show the results in Figure 4.
In total, we marked 49.5% of LLM-generated annotations and 50.5% of reasons as correct (with 9.2% of annotations and 12.5% of reasons additionally marked as partially correct). Reasoning models perform better, with 56.4% of their annotations and 58.3% marked as completely correct. The most accurate annotations on D2T-Eval were those made by Gemini 2.0 and DeepSeek-R1. o3-mini performed well on both tasks, although Propaganda proved to be challenging for all models.
What are the sources of model errors?
Qualitatively, we find that the models often select wrong error categories despite identifying real issues (e.g. labeling contradictory statements as "incoherent"). Models also tend to be overly attentive, flagging noise in the data (quotes or non-article content) as errors, or marking slight numerical variations (such as rounded values) as misleading. All of these cases could be tackled by more descriptive guidelines or additional examples. However, in some cases, the models also misread or misinterpreted the data (e.g., claiming wind speed measurements do not exist when they do), which hints at deeper issues with understanding the data. Incorrect explanations vary from incomplete explanations (addressing only part of a multi-issue span), irrelevant explanations (e.g., counting omissions as sources of errors) to incorrectly flagging subjective statements (e.g., “long-lasting usage”) as factual errors. Occasionally, the model admits that it marked a correct span as an error, such as in “The description of the game’s duration aligns with the data, providing coherent information”. [^14] This typically happens to GPT-4o, even though OpenAI API ensures JSON key ordering so the explanation should have been generated before the annotation (cf. Section 3.2).
How good are human annotations?
Interestingly, the LLM annotations that were marked as correct have only 24% hard character-level overlap (51% soft) with human annotations. This fact led us to analyze the quality of human annotations in D2T-Eval. On a sample of 108 human annotations, we marked 45.3% of the annotations and 54.6% of the explanations correct, which is comparable to the LLM annotations. This suggests that the task is hard even for human annotators, and the quality of annotations from crowdworkers varies, even if they are preselected using a qualification task.
Figure 4: Results of our manual analysis. We analyzed 18 annotations and their explanations for each model and task (216 annotations in total). The color bars show annotations that we classified as correct, partially correct, assigned wrong category, incorrect, and undecidable cases. Detailed results are provided in Tables 22 and 23 in Appendix E.
6 Discussion
Can LLMs substitute human annotators?
The IAA between LLMs and human annotators is generally moderate. However, we also need to take into account a realistic upper bound: the average IAA between human annotators themselves. In that case, the strongest LLMs have already reached the level of skilled crowdworkers in D2T-Eval. As we also showed, not all human annotations are “gold”, as even qualified crowdworkers make similar amounts of errors as LLMs. If we also factor in financial aspects, LLMs provide better flexibility, shorter response times, and lower costs. In light of this, we argue that the advantages of automating span annotation with LLMs outweigh its issues. LLM-based span annotation will also benefit from future increases in LLM capabilities.
Recommendations
To use an LLM as a span annotator, we strongly recommend providing detailed guidelines describing how to handle all edge cases. In contrast, we note that providing specific examples (i.e., a few-shot setup) does not bring consistent improvements, possibly due to the length and complexity of examples for this task. When using LLMs in a new setup, we recommend validating model’s annotations against examples hand-annotated by experts on a sample of the data. Regarding model choice, we recommend choosing a reasoning model as opposed to an equivalent model without reasoning, since the reasoning models provide more reliable annotations at the cost of slightly higher response times and token count.
Limitations
Span annotation as an evaluation method is not suitable for certain NLG evaluation tasks, such as annotating omissions or rating the overall text style. Although we tried to experiment with a representative sample of models, prompts, and tasks, our choice is necessarily constrained by our limited time frame and budget. Our estimates of the upper-bound IAA for each task are difficult to establish and depend on many factors, such as the chosen annotation categories, their inherent ambiguity, the annotation guidelines, or the qualification level of human annotators.
7 Conclusion
We demonstrated that LLMs can serve as span annotators for three span annotation tasks: evaluating data-to-text generation, identifying errors in machine translation, and detecting propaganda in human-written texts. Our experiments show that LLMs achieve moderate agreement with skilled human annotators. The models perform best in D2T-Eval, where they are comparable to qualified crowdworkers. Reasoning models consistently outperform their instruction-tuned counterparts, delivering more accurate annotations and providing more valid explanations for their decisions. Overall, automating span annotation with LLMs offers a promising alternative to human annotation, opening the way for more fine-grained NLG evaluation methods.
Author Contributions
DG and SB first came up with the idea for the project, with SB further coordinating and overseeing the research process. ZK led the experimental design and execution part, including conducting both preliminary and main experiments, organizing the crowdsourcing campaigns, and processing the collected data. Multiple authors (DG, IK, KO, SB, SM, VZ, ZK) participated in the collection of gold data for D2T-Eval. Similarly, multiple authors (IK, KO, OD, OP, PS, SB, ZK) were involved in manual evaluation of the model outputs. DG provided financial resources for the Prolific campaigns. SM and SB provided expertise in preparing annotation guidelines and structuring the Prolific campaigns. Data processing and analysis were handled mainly by ZK, VZ, and PS, with VZ providing extra support with the WMT data. The paper was written by ZK, VZ, IK, PS, OD, and SB.
Acknowledgments
This work was funded by the European Union (ERC, NG-NLG, 101039303) and Charles University project SVV 260 698. It used resources of the LINDAT/CLARIAH-CZ Research Infrastructure (Czech Ministry of Education, Youth, and Sports project No. LM2018101).
We thank David M. Howcroft for his early input and contributions to the research methodologies adopted in this study.
Ethics Statement
The human evaluation study was approved by the internal ethics committee of our institution. Our human annotators were hired over Prolific and paid the platform-recommended wage of 9 GBP/hour (adjusted to slightly higher rates to account for real annotation times). Annotators were pre-selected on the basis of their primary language (English). All annotators were shown detailed instructions and explanation of the data types, data sources, and the purpose of the research (see Appendix B.2 for details). The domains were selected so that they do not contain sensitive or potentially offensive content. We do not collect demographic data about participants.
Appendix A Implementation Details
A.1 Open Models
We run the local models using the ollama framework in 4-bit quantization. Specifically, we use the following local models for span annotations:
We also use the following models for generating texts in D2T-Eval:
For better reproducibility, we set the seed to 42 and the temperature to 0 for the local models. We do not use these parameters for proprietary models as these parameters are generally not supported.
We run the models on a set of GPUs, including NVIDIA H100 NVL (95G), AMD MI210 (64G), and NVIDIA RTX 3090 (24G).
A.2 Proprietary Models
We use the following proprietary model versions:
  • GPT-4o: gpt-4o-2024-11-20
  • Claude 3.7 Sonnet: claude-3-7-sonnet-20250219
  • o3-mini: o3-mini-2025-01-31
  • Gemini 2.0 Flash Thinking: gemini-2.0-flash-thinking-exp-01-21
A.3 Parsing the Output with Reasoning Traces
At the time we performed our experiments, some of the reasoning models either did not support structured output (Gemini 2.0 Flash Thinking) or it would cancel out their ability to perform reasoning (ollama models [^15] https://github.com/ollama/ollama/issues/8529). Therefore, we use the following heuristic: we retrieve the raw answer from the model, strip any parts within the <think></think> tags (if present), and consider the latest valid top-level JSON object as the model’s response.
Appendix B D2T-Eval Details
In this section, we provide more details on the process of collecting data (Section B.1) and annotations (Section B.2) for D2T-Eval.
B.1 Collecting Data
To collect a novel dataset for D2T-Eval that is not subject to data contamination (cf. Section 4.4), we follow the approach of Kasner & Dušek (2024) to collect structured data from publicly available APIs.
Domains
The Quintd tool Kasner & Dušek (2024) enables downloading data for five domains: openweather, gsmarena, ice_hockey, our_world_in_data, and wikidata. We selected openweather and gsmarena as the most suitable for the data collection process. Additionally, we collect the new football domain from RapidAPI - API-Football. [^16] Its structured data contain more detailed information about the game than ice_hockey and the sport itself is also more generally well-known.
Generating model outputs.
To obtain output texts, we prompt two state-of-the-art models, Llama 3.3 Grattafiori et al. (2024) and GPT-4o Hurst et al. (2024), and two smaller models, Gemma 2 Team et al. (2024) and Phi-3.5 Abdin et al. (2024) (cf. Section A.1). We prompt the models in a zero-shot setting, asking them to generate a summary of the given data using approximately five sentences. We provide the prompts in Appendix D.
Note on annotating models’ own outputs.
Two of the models (Llama 3.3 and GPT-4o) are also used later to collect annotations, which can reveal self-preference bias Zheng et al. (2023); Wataoka et al. (2024). Since we do not rank the quality of generated outputs, we do not consider it a flaw but rather a ground for future investigation. Moreover, we hypothesize that since span annotations are more grounded than direct assessment, the self-preference bias will not be as prominent.
B.2 Collecting annotations
Here we describe our process of writing annotation guidelines and selecting crowdworkers in order to ensure high quality of the annotations collected for D2T-Eval.
Annotation guidelines
For the annotation guidelines, we (the authors of this paper) went through an iterative process to establish the annotation guidelines $\mathcal{G}$ and the annotation categories $C$. We started with a preliminary version of the guidelines and annotation categories, drawing inspiration from the guidelines in previous works Kasner & Dušek (2024); Thomson & Reiter (2020). Eventually, we settled on the list of annotation categories described in Appendix C (Table 5).
Our list is an extended and updated version of the error taxonomy of Kasner & Dušek (2024), also covering coherence categories (Incoherent, Repetitive) on top of semantic accuracy.
Additionally, we agreed upon a list of general recommendations for collecting the annotations:
  • The annotator should always annotate the longest continuous span (i.e., the whole fact instead of a single word).
  • The annotator should annotate according to their own knowledge and only the spans that they are sure about.
  • The annotator should ignore subjective statements such as "a lightweight smartphone".
  • The outputs in openweather can contain both precise numbers (e.g. 10.71°C) and the rounded ones (e.g. 11°C) as long as they agree with the data.
We provide these additional recommendations in $\mathcal{P}_{\texttt{base}}$, while we omit them in $\mathcal{P}_{\texttt{noguide}}$ (cf. Figures 6 and 7).
Gold annotations
With the annotation guidelines established, we proceeded to collect our own internally annotated gold data. The purpose of this data is two-fold: (1) to pre-select skilled crowdworkers, and (2) to create a smaller high-quality development set for the model prompting study. Since we also wanted to quantify our own internal IAA, we divided the data into two sets: $\mathcal{D}_{\texttt{dev}}$, containing 84 examples annotated individually by 7 annotators (12 examples per annotator), and $\mathcal{D}_{\texttt{iaa}}$, containing 12 examples annotated commonly by each annotator. Our average IAA on $\mathcal{D}_{\texttt{iaa}}$ was $F_{1}=0.444$ and $\gamma=0.399$.
Qualification task
To gather a pool of skilled annotators, we set up a qualification task on Prolific.com. For the study, we pre-selected workers whose first language is English, with >99% approval rate and more than 100 submissions. The workers were first presented with a detailed tutorial with annotation guidelines and examples of individual errors. After the tutorial, we tested the worker performance on five manually pre-selected examples from $\mathcal{D}_{\texttt{dev}}$. For annotating the data in $\mathcal{D}_{\texttt{test}}$, we selected workers that had $F_{1}$ score higher than 0.5 w.r.t. our ground-truth solution to the qualification task. Of the 230 annotators who participated in the qualification task, 50 annotators (21.7%) passed this baseline. We pay all the annotators an average reward of 9.58 GBP / hour, regardless of the qualification outcome.
Annotating test data
We invited annotators who passed the qualification task to annotate $\mathcal{D}_{\texttt{test}}$. Out of 50 selected participants, 45 participated in the main task (=90% turnover rate). We presented each annotator with a batch of 32 examples: 25 examples from $\mathcal{D}_{\texttt{test}}$ and 7 remaining examples from $\mathcal{D}_{\texttt{dev}}$ (that is, the examples from $\mathcal{D}_{\texttt{dev}}$ that we did not use for the qualification task). All 1200 outputs in $\mathcal{D}_{\texttt{test}}$ were annotated by at least one annotator. Furthermore, 475 outputs (39.6%) were annotated by an additional annotator. We payed all the annotators an average reward of 10.70 GBP / hour. We use examples with two annotators to compute the average IAA for D2T-Eval in Section 5.2.171717For other experiments, we use only the outputs from the first annotator as the reference data.
B.3 Annotation Interface
We implement our annotation interface using factgenie. Figure 5 shows the main annotation interface, including an example from the openweather domain.
Figure 5: Screenshot of the factgenie annotation interface used for collecting span annotations.
Appendix C Annotation Categories
Tables 5, 6 and 7 show an overview of the annotation span categories that we used for our tasks along with their descriptions.
Category Name Description
Contradictory The fact contradicts the data.
Not checkable The fact cannot be verified from the data.
Misleading The fact is technically true, but leaves out important information or otherwise distorts the context.
Incoherent The text uses unnatural phrasing or does not fit the discourse.
Repetitive The fact has been already mentioned earlier in the text.
Other The text is problematic for another reason.
Table 5: Annotation categories for the D2T-Eval task.
Category Name Description
Major An error that disrupts the flow and makes the understandability of text difficult or impossible.
Minor An error that does not disrupt the flow significantly and what the text is trying to say is still understandable.
Table 6: Annotation categories for the MT-Eval task.
Category Name Description
Appeal to Authority Stating that a claim is true simply because a valid authority or expert on the issue said it was true, without any other supporting evidence offered. We consider the special case in which the reference is not an authority or an expert in this technique, altough it is referred to as Testimonial in literature
Appeal to fear-prejudice Seeking to build support for an idea by instilling anxiety and/or panic in the population towards an alternative. In some cases the support is built based on preconceived judgements
Bandwagon Attempting to persuade the target audience to join in and take the course of action because "everyone else is taking the same action"
Black-and-White Fallacy Presenting two alternative options as the only possibilities, when in fact more possibilities exist. As an the extreme case, tell the audience exactly what actions to take, eliminating any other possible choices (Dictatorship)
Causal Oversimplification Assuming a single cause or reason when there are actually multiple causes for an issue. It includes transferring blame to one person or group of people without investigating the complexities of the issue
Doubt Questioning the credibility of someone or something
Exaggeration, Minimisation Either representing something in an excessive manner: making things larger, better, worse (e.g., "the best of the best", "quality guaranteed") or making something seem less important or smaller than it really is (e.g., saying that an insult was just a joke)
Flag-Waving Playing on strong national feeling (or to any group; e.g., race, gender, political preference) to justify or promote an action or idea
Loaded Language Using specific words and phrases with strong emotional implications (either positive or negative) to influence an audience
Name Calling, Labeling Labeling the object of the propaganda campaign as either something the target audience fears, hates, finds undesirable or loves, praises
Obfuscation, Intentional Vagueness, Confusion Using words which are deliberately not clear so that the audience may have its own interpretations. For example when an unclear phrase with multiple definitions is used within the argument and, therefore, it does not support the conclusion
Red Herring Introducing irrelevant material to the issue being discussed, so that everyone“s attention is diverted away from the points made
Reductio ad hitlerum Persuading an audience to disapprove an action or idea by suggesting that the idea is popular with groups hated in contempt by the target audience. It can refer to any person or concept with a negative connotation
Repetition epeating the same message over and over again so that the audience will eventually accept it
Slogans A brief and striking phrase that may include labeling and stereotyping. Slogans tend to act as emotional appeals
Straw Men When an opponent“s proposition is substituted with a similar one which is then refuted in place of the original proposition
Thought-terminating Cliches Words or phrases that discourage critical thought and meaningful discussion about a given topic. They are typically short, generic sentences that offer seemingly simple answers to complex questions or that distract attention away from other lines of thought
Whataboutism A technique that attempts to discredit an opponent“s position by charging them with hypocrisy without directly disproving their argument
Table 7: Annotation categories for the MT-Eval task. The categories are adopted from Da San Martino et al. (2019).
Appendix D Prompts
Figures 6, 7, 8 and 9 show the prompts for the D2T-Eval that we use for the experiments in Section 5.1.
Figure 10 shows the base prompt we used for MT-Eval, and Figure 11 shows the base prompt we used for Propaganda.
Finally, Figure 12 shows the prompt we used for generating the outputs for D2T-Eval.
Your task is to identify errors in the text and classify them.
Output the errors as a JSON object with a single key "annotations". The value of "annotations" is a list in which each object contains fields "reason", "text", and "annotation_type". The value of "reason" is the short sentence justifying the annotation. The value of "text" is the literal value of the identified span (we will later identify the span using string matching). The value of "annotation_type" is an integer index of the error based on the following list:
{categories}
Examples:
- Contradictory: The lowest temperature does not drop below 4°C, but the text mentions 2°C. - Not checkable: The text mentions that "both teams display aggressive play", which cannot be checked from the data.
- Misleading: The tone of the text suggests there are many sensors out of which just a few are listed here. However, according to the data, the device has exactly these four sensors.
- Incoherent: The text states that both teams had "equal chances until the first half ended scoreless." While this is technically true, the expression sounds unnatural for a sport summary.
- Repetitive: The output text unnecessarily re-states information about a smartphone battery that was mentioned earlier.
- Other: Use this as a last resort when you notice something else not covered by the above categories.
Hints:
- Always try to annotate the longest continuous span (i.e., the whole fact instead of a single word).
- Annotate only the spans that you are sure about. If you are not sure about an annotation, skip it.
- Ignore subjective statements: for example "a lightweight smartphone" highly depends on the context: you should not annotate these statements.
- Numerical conventions: For weather forecasts, we accept both precise numbers (e.g. 10.71°C) and the rounded ones (e.g. 11°C) as long as they agree with the data.
- Annotate only according to your own knowledge. If you are considering using an external tool (Google, ChatGPT etc.), just skip that specific fact.
If there is nothing to annotate in the text, "annotations" will be an empty list.
Given the data:
{data} ```
annotate the errors in the corresponding text generated from the data:
{text} ```
Figure 6: The prompt $\mathcal{P}_{\texttt{base}}$ for D2T-Eval.
Your task is to identify errors in the text and classify them.
Output the errors as a JSON object with a single key "annotations". The value of "annotations" is a list in which each object contains fields "reason", "text", and "annotation_type". The value of "reason" is the short sentence justifying the annotation. The value of "text" is the literal value of the identified span (we will later identify the span using string matching). The value of "annotation_type" is an integer index of the error based on the following list:
{categories}
Given the data:
{data} ```
annotate the errors in the corresponding text generated from the data:
{text} ```
Figure 7: The prompt $\mathcal{P}_{\texttt{noguide}}$ for D2T-Eval.
Think about it step-by-step. You should enclose your chain of thoughts between the <think> and </think> tags. Once you are ready, output the JSON object in the required format.
Example:
<think> ... chain of thoughts ... </think> ... JSON object ... ```
Figure 8: The additional text added for $\mathcal{P}_{\texttt{cot}}$.
In order to help you with the task, we provide you with five examples of inputs, outputs and annotations:
Example #1: data:
{data} ```
text:
{text} ```
output:
{annotations} ``` (...)
Figure 9: The additional text added for $\mathcal{P}_{\texttt{5shot}}$.
Your task is to identify errors in the translation and classify them.
Output the errors as a JSON object with a single key "annotations". The value of "annotations" is a list in which each object contains fields "reason", "text", and "annotation_type". The value of "reason" is the short sentence justifying the annotation. The value of "text" is the literal value of the identified span (we will later identify the span using string matching). The value of "annotation_type" is an integer index of the error based on the following list:
{categories}
Error spans can include parts of the words, whole words, or multi-word phrases. Hint: errors are usually accuracy-related (addition, mistranslation, omission, untranslated text), fluency-related (character encoding, grammar, inconsistency, punctuation, register, spelling), style-related (awkward), terminology (inappropriate for context, inconsistent use).
Make sure that the annotations are not overlapping. If there is nothing to annotate in the text, "annotations" will be an empty list.
Given the source:
{source} ```
annotate its translation:
{text} ```
Figure 10: The prompt $\mathcal{P}_{\texttt{base}}$ for MT-Eval.
Your task is to identify spans of text that employ propaganda techniques.
Output the errors as a JSON object with a single key "annotations". The value of "annotations" is a list in which each object contains fields "reason", "text", and "annotation_type". The value of "reason" is the short sentence justifying the annotation. The value of "text" is the literal value of the identified span (we will later identify the span using string matching). The value of "annotation_type" is an integer index of the error based on the following list:
{categories}
Now annotate the following text:
{text} ```
Figure 11: The prompt $\mathcal{P}_{\texttt{base}}$ for Propaganda.
Given the structured summary of a football game:
{data} ``` Generate a match summary using approximately five sentences. The summary should sound natural, reporting on the important moments of the game. Avoid subjective statements, keep the tone of the summary neutral. Do not fabricate any facts that are not explicitly stated in the data.
Figure 12: The prompt used for generating outputs in the football domain for D2T-Eval. The prompts for the other domains are analogical. For more robust parsing, we initialize the model response with ’Sure, here is the summary: "’ .
Appendix E Results
In this section, we provide detailed results of our experiments.
The results for D2T-Eval are included in Sections E.1, E.2, E.3 and E.4:
The results for MT-Eval are included in Sections E.5 and E.6:
The results for Propaganda are included in Section E.7: Tables 20 and 21.
The confusion matrices are included in Section E.8: Figures 13, 14 and 15.
The manual evaluation results are included in Section E.9: Tables 22 and 23.
E.1 D2T-Eval – Main Results
Model $\rho$ Precision Recall F1 $\gamma$ $S_{\emptyset}$
Hard Soft Hard Soft Hard Soft $\Delta$
Llama 3.3 0.307 0.176 0.365 0.187 0.388 0.181 0.377 0.196 0.109 0.418
GPT-4o 0.346 0.233 0.391 0.184 0.308 0.206 0.345 0.139 0.130 0.429
Claude 3.7 0.512 0.294 0.442 0.304 0.457 0.299 0.449 0.150 0.203 0.592
DeepS. R1 0.453 0.317 0.532 0.185 0.310 0.233 0.392 0.159 0.185 0.645
o3-mini 0.505 0.392 0.542 0.285 0.395 0.330 0.457 0.127 0.273 0.637
Gem. 2-FT 0.458 0.293 0.488 0.263 0.438 0.277 0.462 0.185 0.209 0.612
Table 8: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on D2T-Eval. See Figure 2 for visualizaton of this table.
Annotator Ann. Ann/Ex Ex. w/o annChars/Ann
Human 2981 2.5 28.8 50.3
Llama 3.3 3214 2.7 7.4 65.5
GPT-4o 2284 1.9 4.8 66.3
Claude 3.7 2865 2.4 22.5 57.2
DeepS. R1 1387 1.2 44.2 56.8
o3-mini 1836 1.5 35.6 58.0
Gem. 2-FT 2517 2.1 28.9 54.3
Table 9: Statistics of models and human annotators using $\mathcal{P}_{\texttt{base}}$ on D2T-Eval. Ann=# of annotations, Ann/Ex=annotations per example. w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
E.2 D2T-Evalfootball domain
Model $\rho$ Precision Recall F1 $\gamma$ $S_{\emptyset}$
Hard Soft Hard Soft Hard Soft $\Delta$
Llama 3.3 0.512 0.288 0.484 0.296 0.498 0.292 0.491 0.199 0.197 0.455
GPT-4o 0.422 0.347 0.502 0.268 0.388 0.302 0.438 0.136 0.190 0.427
Claude 3.7 0.612 0.360 0.511 0.391 0.555 0.375 0.532 0.157 0.254 0.587
DeepS. R1 0.561 0.431 0.670 0.275 0.428 0.336 0.522 0.186 0.276 0.659
o3-mini 0.610 0.470 0.608 0.384 0.496 0.422 0.547 0.125 0.328 0.645
Gem. 2-FT 0.588 0.419 0.617 0.347 0.512 0.379 0.560 0.181 0.307 0.647
Table 10: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the football domain of D2T-Eval.
Annotator Ann. Ann/Ex Ex. w/o annChars/Ann
Human 1053 2.6 26.8 54.1
Llama 3.3 977 2.4 6.2 71.5
GPT-4o 761 1.9 2.8 72.8
Claude 3.7 1051 2.6 13.8 63.1
DeepS. R1 486 1.2 41.2 67.1
o3-mini 642 1.6 30.5 70.8
Gem. 2-FT 755 1.9 29 61.3
Table 11: Statistics of models and human annotators using $\mathcal{P}_{\texttt{base}}$ on the football domain of D2T-Eval. Ann=# of annotations, Ann/Ex=annotations per example. w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
E.3 D2T-Evalgsmarena domain
Model $\rho$ Precision Recall F1 $\gamma$ $S_{\emptyset}$
Hard Soft Hard Soft Hard Soft $\Delta$
Llama 3.3 0.165 0.107 0.247 0.122 0.282 0.114 0.263 0.149 0.077 0.456
GPT-4o 0.127 0.107 0.269 0.082 0.205 0.093 0.233 0.140 0.062 0.453
Claude 3.7 0.214 0.176 0.348 0.152 0.300 0.163 0.322 0.159 0.140 0.656
DeepS. R1 0.223 0.199 0.405 0.089 0.180 0.123 0.249 0.126 0.160 0.686
o3-mini 0.172 0.225 0.473 0.115 0.241 0.152 0.320 0.168 0.192 0.689
Gem. 2-FT 0.234 0.218 0.426 0.128 0.249 0.161 0.314 0.153 0.161 0.675
Table 12: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the gsmarena domain of D2T-Eval.
Annotator Ann. Ann/Ex Ex. w/o annChars/Ann
Human 809 2.0 40.2 42.5
Llama 3.3 890 2.2 15 73.9
GPT-4o 670 1.7 10.5 59.2
Claude 3.7 584 1.5 47.2 48.9
DeepS. R1 208 0.5 69.2 53.8
o3-mini 272 0.7 66.2 49.1
Gem. 2-FT 419 1.1 54.2 45.1
Table 13: Statistics of models and human annotators using $\mathcal{P}_{\texttt{base}}$ on the gsmarena domain of D2T-Eval. Ann=# of annotations, Ann/Ex=annotations per example. w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
E.4 D2T-Evalopenweather domain
Model $\rho$ Precision Recall F1 $\gamma$ $S_{\emptyset}$
Hard Soft Hard Soft Hard Soft $\Delta$
Llama 3.3 0.209 0.112 0.327 0.118 0.344 0.115 0.335 0.220 0.053 0.272
GPT-4o 0.404 0.196 0.354 0.160 0.289 0.176 0.318 0.142 0.122 0.377
Claude 3.7 0.550 0.274 0.411 0.292 0.437 0.283 0.424 0.141 0.189 0.429
DeepS. R1 0.454 0.241 0.436 0.143 0.258 0.180 0.324 0.144 0.124 0.533
o3-mini 0.552 0.362 0.496 0.267 0.366 0.307 0.422 0.115 0.259 0.481
Gem. 2-FT 0.458 0.226 0.415 0.249 0.458 0.237 0.435 0.198 0.155 0.368
Table 14: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the openweather domain of D2T-Eval.
Annotator Ann. Ann/Ex Ex. w/o annChars/Ann
Human 1119 2.8 19.2 52.2
Llama 3.3 1347 3.4 1 55.7
GPT-4o 853 2.1 1.2 66.2
Claude 3.7 1230 3.1 6.5 56.1
DeepS. R1 693 1.7 22.2 50.5
o3-mini 922 2.3 10 51.8
Gem. 2-FT 1343 3.4 3.5 53.2
Table 15: Statistics of models and human annotators using $\mathcal{P}_{\texttt{base}}$ on the openweather domain of D2T-Eval. Ann=# of annotations, Ann/Ex=annotations per example. w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
E.5 MT-Eval – Main Results
Model $\rho$ Precision Recall F1 $\gamma$ $S_{\emptyset}$
Hard Soft Hard Soft Hard Soft $\Delta$
Llama 3.3 0.182 0.121 0.200 0.229 0.378 0.155 0.257 0.102 0.014 0.392
GPT-4o 0.158 0.141 0.240 0.195 0.327 0.156 0.266 0.110 0.076 0.428
Claude 3.7 0.301 0.226 0.325 0.335 0.469 0.262 0.373 0.111 0.131 0.628
DeepS. R1 0.177 0.169 0.268 0.183 0.280 0.168 0.262 0.094 0.058 0.631
o3-mini 0.197 0.169 0.291 0.161 0.275 0.160 0.275 0.115 0.100 0.646
Gem. 2-FT 0.257 0.184 0.312 0.180 0.339 0.173 0.304 0.130 0.066 0.710
Table 16: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the MT-Eval – average across languages.
Annotator Ann. Ann/Ex Ex. w/o annChars/Ann
Human 2090 0.7 66.0 14.5
Llama 3.3 6361 2.3 6.2 17.4
GPT-4o 4866 1.7 7.0 15.9
Claude 3.7 3782 1.4 30.6 15.9
DeepS. R1 2586 0.9 36.3 15.1
o3-mini 3039 1.1 35.8 13.8
Gem. 2-FT 2181 0.8 50.0 15.2
Table 17: Statistics of models and human annotators using $\mathcal{P}_{\texttt{base}}$ on the MT-Eval – average across languages. Ann=# of annotations, Ann/Ex=annotations per example. w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
E.6 MT-Eval – By Language
Model $\rho$ Precision Recall F1 $\gamma$ $S_{\emptyset}$
Hard Soft Hard Soft Hard Soft $\Delta$
en-cs 0.303 0.144 0.268 0.180 0.326 0.156 0.286 0.130 0.084 0.582
en-es 0.171 0.161 0.243 0.236 0.362 0.190 0.288 0.098 0.080 0.631
en-hi 0.170 0.173 0.265 0.208 0.327 0.173 0.269 0.096 -0.0 0.552
en-is 0.347 0.136 0.246 0.187 0.361 0.145 0.269 0.124 0.108 0.493
en-ja 0.127 0.193 0.302 0.249 0.363 0.209 0.318 0.109 0.063 0.569
en-ru 0.225 0.178 0.256 0.273 0.386 0.208 0.298 0.090 0.162 0.588
en-uk 0.192 0.166 0.254 0.214 0.339 0.184 0.286 0.102 0.031 0.542
en-zh 0.163 0.196 0.346 0.163 0.294 0.169 0.302 0.133 0.075 0.623
Table 18: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the MT-Eval separately for each language (average across models).
Language Annotator Ann. Ann/Ex Ex. w/o annChars/Ann
en-cs Model avg. 600 1.4 27.0 16.6
Human 399 0.7 66.1 13.0
en-es Model avg. 417 1.1 38.9 18.8
Human 248 0.6 70.3 10.3
en-hi Model avg. 396 1.3 26.2 19.0
Human 222 0.5 71.2 10.7
en-is Model avg. 563 1.9 14.3 15.7
Human 752 2.5 18.3 16.6
en-ja Model avg. 471 1.3 28.7 11.1
Human 118 0.2 87.5 14.8
en-ru Model avg. 500 1.3 25.9 18.2
Human 287 0.7 58.7 19.4
en-uk Model avg. 436 1.5 25.4 17.8
Human 208 0.7 64.3 12.3
en-zh Model avg. 420 1.2 34.6 7.2
Human 171 0.2 85.1 6.6
Table 19: Statistics of models (averaged) and human annotators using $\mathcal{P}_{\texttt{base}}$ on the MT-Eval separately for each language. Ann=# of annotations, Ann/Ex=annotations per example. Ex. w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
E.7 Propaganda – Main Results
Model $\rho$ Precision Recall F1 $\gamma$ $S_{\emptyset}$
Hard Soft Hard Soft Hard Soft $\Delta$
Llama 3.3 0.336 0.070 0.243 0.063 0.219 0.066 0.230 0.164 0.092 0.343
GPT-4o 0.344 0.095 0.293 0.038 0.115 0.054 0.166 0.112 0.066 0.234
Claude 3.7 0.460 0.110 0.274 0.096 0.239 0.103 0.255 0.152 0.155 0.113
DeepS. R1 0.354 0.083 0.246 0.062 0.182 0.071 0.209 0.138 0.091 0.476
o3-mini 0.418 0.152 0.411 0.066 0.179 0.092 0.249 0.157 0.154 0.517
Gem. 2-FT 0.560 0.106 0.268 0.190 0.477 0.136 0.343 0.207 0.202 0.493
Table 20: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the Propaganda.
Annotator Ann. Ann/Ex Ex. w/o annChars/Ann
Human 1439 14.2 4.0 40.2
Llama 3.3 574 5.7 3.0 92.0
GPT-4o 246 2.4 8.9 91.1
Claude 3.7 803 8.0 7.9 58.5
DeepS. R1 459 4.5 9.9 89.3
o3-mini 376 3.7 10.9 65.3
Gem. 2-FT 1864 18.5 3.0 54.1
Table 21: Statistics of models and human annotators using $\mathcal{P}_{\texttt{base}}$ on the Propaganda. Ann=# of annotations, Ann/Ex=annotations per example. w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
E.8 Confusion Matrices
(a) Llama 3.3
(b) DeepSeek-R1
(c) GPT-4o
(d) o3-mini
(e) Gemini 2.0 Flash Thinking
(f) Claude 3.7 Sonnet
Figure 13: Confusion matrices comparing human annotations (rows) with model predictions (columns) for D2T-Eval. Categories: Contradictory, Not checkable, Misleading, Incoherent, Repetitive, Other (cf. Table 5). For the model average, see Figure 3.
(a) Llama 3.3
(b) DeepSeek-R1
(c) GPT-4o
(d) o3-mini
(e) Gemini 2.0 Flash Thinking
(f) Claude 3.7 Sonnet
Figure 14: Confusion matrices comparing human annotations (rows) with model predictions (columns) for MT-Eval. Categories: Major and Minor errors (cf. Table 6).
Figure 15: Confusion matrix of category counts averaged across models for Propaganda (see Table 7 for the description of categories).
E.9 Manual Analysis
Annotation Reason
Model C P W I U C P W I U
Llama 3.3 7 1 2 8 0 6 0 1 10 1
GPT-4o 5 2 1 10 0 6 1 0 11 0
Claude 3.7 7 2 3 6 0 9 2 0 7 0
DeepSeek 11 2 4 1 0 12 3 0 3 0
o3-mini 10 3 2 0 3 8 7 0 0 3
Gemini 2 F-T 12 4 1 1 0 12 5 1 0 0
Total 52 14 13 26 3 53 18 2 31 4
Table 22: Manual evaluation results for D2T-Eval domain. Categories for annotation and reason: C=Correct, P=Partially correct, W=Wrong category, I=Incorrect, U=Undecidable.
Annotation Reason
Model C P W I U C P W I U
Llama 3.3 9 1 2 5 1 8 2 1 6 1
GPT-4o 9 0 3 6 0 8 2 1 7 0
Claude 3.7 9 2 3 4 0 9 3 3 3 0
DeepSeek 6 0 4 8 0 7 0 3 8 0
o3-mini 15 0 0 2 1 15 0 0 2 1
Gemini 2 F-T 7 3 2 6 0 9 2 1 6 0
Total 55 6 14 31 2 56 9 9 32 2
Table 23: Manual evaluation results for Propaganda domain. Categories for annotation and reason: C=Correct, P=Partially correct, W=Wrong category, I=Incorrect, U=Undecidable.
Appendix F Related Work (extended)
This section provides more discussion on related work following Section 2.
Due to the complexity and often open-ended nature of generated text, NLG evaluation is inherently difficult (Celikyilmaz et al., 2020). Machine translation (MT) has historically played an important role in the development of both automated evaluation metrics and human annotation protocols. Although evaluation practices are less standardized in other tasks, human evaluation typically involves ranking outputs or assigning numeric scores for specific aspects such as fluency or faithfulness (Novikova et al., 2018; van der Lee et al., 2021; Kocmi et al., 2022).
Error Span Annotation.
As simple numeric scores can be subjective and unreliable, the field is shifting towards annotating erroneous spans. The MT community has adopted a detailed error taxonomy, known as MQM (Lommel et al., 2014; Mariana, 2014, Multidimensional Quality Metrics). In MQM, each erroneous span in the translation has an assigned category, such as accuracy/terminology and severity, such as minor/major/critical. The final translation score is defined as a weighted sum of the error spans based on the error categories and severities. However, this leads to many ties and is problematic when judging longer translations. In addition, this protocol requires a very expensive expert human labor. For these reasons, Kocmi et al. (2024c), based on prior work by Popovic (2020), introduce ESA (Error Span Annotation) for machine translation evaluation. This protocol drops the error categorization, which is primarily useful for diagnosis and not evaluation, simplifies severities, and has the annotators also assign the final score after they have done the span annotations. To make large-scale annotation more efficient, Zouhar et al. (2025) use a quality estimation system that pre-highlights potential error spans in the translation. The annotators then “post-edit” the error annotations and assign the final score.
In data-to-text, Thomson & Reiter (2020) introduce an evaluation protocol for high-quality annotation of semantic accuracy errors in generated basketball reports. Similarly to MQM, such an approach involves significant costs and time requirements, as the authors report. Beyond NLG evaluation, span annotation has also been applied in other areas, such as fine-grained propaganda detection. Da San Martino et al. (2019) build an annotated corpus with span annotations for 18 different propaganda techniques, followed by Hasanain et al. (2024) and Ramponi et al. (2025) who apply similar approach for propaganda detection in Arabic and Italian, respectively.
Automated Metrics.
Although human evaluation is considered the gold standard for assessing quality in NLG tasks, its substantial demands in terms of time and cost make it less scalable and practical. This has led to an overreliance on traditional automatic metrics (Schmidtová et al., 2024), which typically assess text quality by measuring lexical overlap with human-written reference texts (e.g., Papineni et al., 2002; Lin, 2004). Although efficient and scalable, these metrics focus on surface-level properties and are unable to measure fine-grained aspects of text quality (Gehrmann et al., 2023). In addition, traditional automatic metrics have been shown to generally correlate weakly with human judgements (Novikova et al., 2017; Reiter, 2018). Model-based metrics (Rei et al., 2020; Yuan et al., 2021; Zhong et al., 2022) have addressed these issues to some extent by fine-tuning pre-trained models to evaluate text beyond lexical overlap. While these metrics generally correlate better with human judgments compared to overlap-based metrics, they still suffer from several limitations. Specifically, they provide limited explainability by providing only numeric scores, require large amounts of labeled training data, and are often limited to evaluating specific tasks or aspects (e.g., Kryscinski et al., 2020; Mehri & Eskénazi, 2020).
To automate error span annotation in data-to-text, Kasner et al. (2021) introduce a combination of a rule-based NLG system with a neural retriever model and fine-tuned RoBERTa (Liu et al., 2019). Goyal et al. (2022) propose a protocol for fine-grained annotation of coherence errors in narrative summarization, applying it to both human and automatic span annotations. In MT evaluation, xCOMET (Guerreiro et al., 2024) is a recent model-based metric specialized in error span annotation.
LLM-as-a-judge.
With the emerging LLM-as-a-judge paradigm (Gu et al., 2024), LLMs have been applied as evaluators across diverse tasks. Although their generative capabilities initially have been applied mainly to simple numeric scoring Liu et al. (2023); Sottana et al. (2023), other approaches include generating textual feedback before their final decision (Li et al., 2024; Kim et al., 2024a; b).
Relatively few works have addressed evaluation as error span annotation with LLMs. To obtain error spans for translation evaluation, GEMBA-MQM (Kocmi & Federmann, 2023) promps an LLM with annotation guidelines based on MQM evaluation protocol. InstructScore (Xu et al., 2023) and TigerScore (Jiang et al., 2024a) fine-tune smaller LLMs on synthetic data generated by GPT-4 to generate error spans. Although computationally more efficient than prompting larger LLMs, these two metrics either require reference texts, or are limited to predefined evaluation aspects. Kartác et al. (2025) apply an ensemble of open-weight LLMs as evaluators to a variety of tasks and aspects, generating both numeric scores and error spans with detailed explanations. To address potential data contamination, Kasner & Dušek (2024) collect novel structured data and model outputs in five domains, and collect span annotations for semantic accuracy errors with both crowdworkers and GPT-4.
Goel et al. (2023) apply LLMs for medical text annotation, identifying and extracting spans containing specific named entities.
Appendix G Examples
In Tables 26, 25 and 24, we show examples of the annotated outputs for our tasks.
In Table 27, we show an example of a reasoning trace of Llama 3.3 prompted with $\mathcal{P}_{\texttt{cot}}$.
In Tables 28 and 29, we show the examples of reasoning traces of DeepSeek-R1.
Annotated by Llama 3 When the left made Linda Sarsour into its role model, it climbed into bed withfear the worst of the worstlabelling. The father of a missing 4-year-old Georgia boy was training children at a filthy New Mexico compoundloaded to commit school shootings, prosecutors alleged in court documents Wednesday.
Annotated by DeepSeek r1 When the left made Linda Sarsour into its role model, it climbed into bed with the worst of the worst.loaded The father of a missing 4-year-old Georgia boy was training children at a filthy New Mexico compound to commit school shootings, prosecutors alleged in court documents Wednesday.exag.
Annotated by o3-mini When the left made Linda Sarsour into its role model, it climbed into bed with the worst of the worst.loaded The father of a missing 4-year-old Georgia boy was training children at a filthy New Mexico compoundloaded to commit school shootings, prosecutors alleged in court documents Wednesday.
Annotated by human When the left made Linda Sarsour into its role modellabelling, it climbed into bedloaded with the worst of the worst.exag. The father of a missing 4-year-old Georgia boy was training children at a filthy New Mexico compoundlabelling to commit school shootings, prosecutors alleged in court documents Wednesday.
Table 24: Two examples for span annotation of Propaganda outputs with appeal-to-fearfear, name-calling-labellinglabelling, loaded-languageloaded, and exaggerationexag. as span categories.
Source Stealthily I made my way towards the back of the store, murmurs and talking heard beyond a metal grate which took a few yanks to tear off. Climbing into the vent was the safest and quietest option for me right now; climbing my way through, I’d pick up on conversations and talks between strangely dressed men. I noticed Al inside, chained. Figures he’d get caught eventually.
Annotated output by Llama3.3 Plížil jsem semajor ke zadní části obchodu, šepot a mluvení se ozývaly za kovovým mřížemiminor, které se podařilo po několika pokusech strhnout. Vlézt do větrací šachty byla pro mě v tuto chvíli nejbezpečnější a nejtišší možnost; když jsem se tam dostal, naslouchal jsem rozhovorům a rozhovorům mezi podivně oblečenými mužiminor. Všiml jsem si Ala uvnitř, v řetězech. Je jasné, že se nakonec chytil.major
Annotated output by DeepSeek r1 Plížil jsem se ke zadní části obchodu, šepot a mluvení se ozývaly za kovovým mřížemi, které se podařilo po několika pokusech strhnout. Vlézt do větrací šachty byla pro mě v tuto chvíli nejbezpečnější a nejtišší možnost; když jsem se tam dostal, naslouchal jsem rozhovorům a rozhovorůmminor mezi podivně oblečenými muži. Všiml jsem si Ala uvnitř, v řetězech. Je jasné, že se nakonec chytil.minor
Annotated output by o3-mini Plížil jsem se ke zadní části obchodu, šepot a mluvení se ozývaly za kovovým mřížemiminor, které se podařilo po několika pokusech strhnout. Vlézt do větrací šachty byla pro mě v tuto chvíli nejbezpečnější a nejtišší možnost; když jsem se tam dostal, naslouchal jsem rozhovorům a rozhovorůmminor mezi podivně oblečenými muži. Všiml jsem si Ala uvnitř, v řetězech. Je jasné, že se nakonec chytil.major
Annotated output by human Plížil jsem se ke zadní části obchodu, šepot a mluvení se ozývaly za kovovým mřížemi, které se podařilo po několika pokusech strhnout. Vlézt do větrací šachty byla pro mě v tuto chvíli nejbezpečnější a nejtišší možnost; když jsem se tam dostal, naslouchal jsem rozhovorům a rozhovorům mezi podivně oblečenými muži. Všiml jsem si Ala uvnitř, v řetězech. Je jasné, že se nakonec chytil.minor
Source “It’s your birthday soon, isn’t it?” Ivory asked, remembering that the princess’ birthday was in a few days. Kari nodded, a sad glint in her light blue eyes.
Annotated output by Llama 3 “¿Es su cumpleaños pronto, no lo es?”major Ivory le preguntó,minor recordando que el cumpleaños de la princesa era en unos días. Kari nodó, un deslumbramiento tristemajor en sus ojos azules claros.
Annotated output by DeepSeek r1 “¿Es su cumpleaños pronto, no lo es?”major Ivory le preguntó, recordando que el cumpleaños de la princesa era en unos días. Kari nodó, un deslumbramientominor triste en sus ojos azules claros.
Annotated output by o3-mini “¿Es su cumpleaños pronto, no lo es?”minor Ivory le preguntó, recordando que el cumpleaños de la princesa era en unos días. Kari nodó, un deslumbramiento tristemajor en sus ojos azules claros.
Annotated output by human “¿Es su cumpleaños pronto, no lo es?”minor Ivory le preguntó, recordando que el cumpleaños de la princesa era en unos días. Kari nodómajor, un deslumbramiento triste en sus ojos azules claros.
[…]
Annotated by Llama 3 Borsec will experience a period of overcast skies with light snow expected for the next few days.not check. Temperatures are expected to remain relatively low, hovering around -7°C to -10°C.not check. There’s a chance of scattered clouds and occasional clear periods throughout the week. Winds will be moderate, ranging from 1-3 m/s.mislead. Overall, Borsec is likely to experience mostly cloudy conditions with some brief periods of clear skies.
Annotated by DeepSeek r1 Borsec will experience a period of overcast skiescont. with light snow expected for the next few days. Temperatures are expected to remain relatively low, hovering around -7°C to -10°C.mislead. There’s a chance of scattered clouds and occasional clear periods throughout the week. Winds will be moderate, ranging from 1-3 m/s. Overall, Borsec is likely to experience mostly cloudy conditions with some brief periods of clear skies.
Annotated by o3-mini Borsec will experience a period of overcast skies with light snow expected for the next few days. Temperatures are expected to remain relatively low, hovering around -7°C to -10°C.mislead. There’s a chance of scattered clouds and occasional clear periods throughout the week. Winds will be moderate, ranging from 1-3 m/s. Overall, Borsec is likely to experience mostly cloudy conditions with some brief periods of clear skies.
Annotated by human Borsec will experience a period of overcast skies with light snow expected for the next few days. Temperatures are expected to remain relatively low, hovering around -7°C to -10°C.cont. There’s a chance of scattered clouds and occasional clear periods throughout the week. Winds will be moderate,cont. ranging from 1-3 m/s. Overall, Borsec is likely to experience mostly cloudy conditions with some brief periods of clear skies.
Table 26: Example for span annotation of D2T-Eval in the weather domain with contradictorycont, misleadingmislead., and not-checkeablenot check. as error span categories. In the annotation interface, the weather panel is interactive and hovering over specific time point shows the temperature and wind conditions.
Annotated by human annotator The Ituano team won 4-3 against Vila Nova in a match played on November 17th, 2024. Ituano scored four goals throughout the game, with Leozinho and Salatiel contributing to their victory.mislead. Vila Nova fought back with three goals from Luciano Naninho, Alesson, and Jemmes, but ultimately fell short.
Annotated by Llama 3 with $\mathcal{P}_{\texttt{base}}$ The Ituano team won 4-3 against Vila Nova in a match played on November 17th, 2024. Ituano scored four goals throughout the game, with Leozinho and Salatiel contributing to their victory.mislead. Vila Nova fought back with three goals from Luciano Naninho, Alesson, and Jemmesmislead., but ultimately fell short.
Annotated by Llama 3 with $\mathcal{P}_{\texttt{cot}}$ The Ituano team won 4-3 against Vila Nova in a match played on November 17th, 2024. Ituano scored four goals throughout the game, with Leozinho and Salatiel contributing to their victory. Vila Nova fought back with three goals from Luciano Naninho, Alesson, and Jemmes, but ultimately fell short.
Reasoning trace of Llama 3 with $\mathcal{P}_{\texttt{cot}}$ (emphasis added) To identify errors in the text, we need to compare it with the provided data. The data contains information about a football match between Ituano and Vila Nova, including the score, goals, and events during the game. First, let’s check if the score mentioned in the text matches the data. According to the data, the score is indeed 4-3 in favor of Ituano, so this part of the text is correct. Next, we look at the goals scored by each team. The text mentions that Ituano scored four goals and attributes contributions to their victory to Leozinho and Salatiel. From the data, we see that Ituano’s goals were scored by José Aldo, Leozinho, Salatiel, and Vinícius Paiva. While Leozinho and Salatiel are correctly mentioned as contributors, the text does not mention José Aldo and Vinícius Paiva, who also scored for Ituano. However, this omission does not necessarily make the statement incorrect or misleading since it doesn’t claim to list all scorers. The text then mentions that Vila Nova scored three goals from Luciano Naninho, Alesson, and Jemmes. According to the data, these players indeed scored for Vila Nova, with their goals being part of the events listed. Therefore, this information is accurate based on the provided data. Given the information in the text and comparing it with the data, there doesn’t seem to be any direct contradictions (Contradictory), unverifiable claims (Not checkable), misleading statements (Misleading), incoherent phrasing (Incoherent), repetitive information (Repetitive), or other issues that would require annotation. Since the text accurately reflects the data provided without introducing any errors or problematic statements, there are no annotations to make.
Table 27: Example of a reasoning trace of Llama 3.3 prompted with $\mathcal{P}_{\texttt{cot}}$ in the D2T-Eval football domain. The model prompted with $\mathcal{P}_{\texttt{cot}}$ dismisses the annotation of the sentence “with Leozinho” that the model prompted with $\mathcal{P}_{\texttt{base}}$ (and the human annotator) annotate as misleadingmislead.. The screenshot of the data input is truncated.
Annotated by DeepSeek-R1 The Posh Equal Plus X700 is a 7-inch smartphone with an IPS LCD display and a resolution of 600 x 1024 pixels. It features a 5MP main camera with LED flash and supports video recording. The device runs on Android 6.0 Marshmallow and has a Qualcomm MT8321 chipsetcont., along with 1GB RAM and 16GB internal storage. The Posh Equal Plus X700 is powered by a Li-Po battery with a capacity of 2650 mAh.
Reasoning trace of DeepSeek-R1 (truncated) Alright, I need to identify errors in the given text based on the provided data. Let me go through each sentence one by one. First sentence: "The Posh Equal Plus X700 is a 7-inch smartphone with an IPS LCD display and a resolution of 600 x 1024 pixels." Checking the data, under ’Display’, it says size is 7.0 inches and resolution 600x1024. So that’s correct. Next: "It features a 5MP main camera with LED flash and supports video recording." Data shows Main Camera has 5 MP and LED flash, Video is ’Yes’. Correct again. Then: "The device runs on Android 6.0 Marshmallow…" Wait, in the data under Platform, OS is Android 6.0 (Marshmallow), correct. But chipset is Mediatek MT8321, not Qualcomm. The text says Qualcomm MT8321, which is wrong. So that’s an error. Next part: "…and has a Qualcomm MT8321 chipset, along with 1GB RAM and 16GB internal storage." As mentioned, chipset is incorrect. Also, Memory section shows Internal as 16GB 1GB RAM, so that part is correct. Last sentence: "The Posh Equal Plus X700 is powered by a Li-Po battery with a capacity of 2650 mAh." Battery details match exactly. So no issue here. So the only error is the chipset being incorrectly stated as Qualcomm instead of Mediatek. This contradicts the data, so it’s a Contradictory error.
Table 28: The reasoning trace of DeepSeek-R1 prompted with $\mathcal{P}_{\texttt{base}}$ in D2T-Eval gsmarena domain with a contradictorycont annotation. The visualization of the input data is truncated, the relevant field is highlighted.
Input text annotated by DeepSeek-R1 (truncated) Dallas Chief Says She Can’t Fire Amber Guyger for Killing Botham Jean - Dept Rules Imply Otherwiseb-w-fall. Dallas, TX — As citizens call for justice in the case of Dallas police officer Amber Guyger — the cop who killed the innocentBotham Jean in his own home — the controversy continues to grow. In a town hall meeting this week, citizens showed that they still have a lot of questions for the Dallas Police Department in regards to the cold-bloodedkilling of Botham Jean. On top of questions as to why it took several days to arrest the cop — despite the fact that Guyger admitted to it — were questions like why this person who has proven to be a threat to societyhas yet to be fired from the police force. During the meeting, as Chief Renee Hall dodged several questions over the details of the arrest, she did go out of her way to answer the question of why Guyger is still collecting a paycheck from their department. The short of it, according to Hall, is she can’t fire her.
[…2,849 characters skipped …]
Reasoning trace of DeepSeek-R1 (truncated) Okay, so I need to analyze the given text and identify spans that use propaganda techniques. Let me go through each paragraph step by step. Starting with the title: "Dallas Chief Says She Can’t Fire Amber Guyger for Killing Botham Jean - Dept Rules Imply Otherwise." Hmm, this seems to present two opposing viewpoints without a middle ground. It’s implying either she can’t fire her or the dept rules say otherwise, which might be a Black-and-White Fallacy. In the first paragraph, it mentions citizens calling for justice and the controversy growing. Not sure about propaganda here yet. The second paragraph talks about a town hall meeting with questions about the arrest delay and why Guyger hasn’t been fired. Still more informative than propagandistic. Moving on, Chief Hall dodges questions but explains she can’t fire Guyger due to laws. Then experts say this isn’t true. This sets up a conflict where the authority (Chief) is contradicted by other experts, which might relate to Appeal_to_Authority if the article is using the experts’ opinions to discredit the Chief without evidence.
[…1,385 characters skipped …]
So, I think several propaganda techniques are present: Appeal_to_Authority (using experts), Black-and-White Fallacy in the title, Exaggeration/Minimisation with other officers’ cases, Doubt when questioning Hall’s actions, and Loaded_Language throughout.
Table 29: Excerpt from the reasoning trace of DeepSeek-R1 prompted with $\mathcal{P}_{\texttt{base}}$ in Propaganda. Both the input text (containing the annotated Black-and-White Fallacyb-w-fall.) and the reasoning trace are truncated.